# PAT: Accelerating LLM Decoding via Prefix-Aware Attention with Resource Efficient Multi-Tile Kernel

### v1

@@ 11. Conclusion
In this work, we present PAT, a prefix-aware attention kernel implementation for LLM decoding. PAT reduces redundant global memory access and improves bandwidth utilization through a pack–forward–merge paradigm. The design incorporates a heuristic pack scheduler, a resource-efficient multi-tile kernel, and forward-stage optimizations that mitigate execution bubbles. We implement PAT as an off-the-shelf plugin to vLLM and demonstrate that it reduces the attention latency by 67.4% on average under synthetic workloads and reduces the TPOT by up to 83.4% under real-world workloads compared with existing works.
Appendix A Extra End-to-end Results
Figure 13. End-to-end performance of PAT and the baselines under two workloads: Llama-3-8B with toolagent and Qwen3-8B with conversation.:
Appendix B Overhead Analysis
The primary overhead introduced by PAT comes from the pack scheduler, which packs decode batches into CTAs at runtime. To mitigate its overhead, as described in §5.1, we design a lazy update mechanism that reduces the triggering frequency of the scheduler without affecting correctness, and overlaps its execution with the serving system’s preprocessing stage. To evaluate the overhead, we measure both the scheduling latency and the preprocessing latency under the toolagent and conversation traces at a request rate of 8 and 5 req/s. As shown in Figure 14, the average scheduling latency is lower than the preprocessing latency by 81.6% and 88.8%, respectively. Therefore, when running in an asynchronous CPU thread, the pack scheduler will not introduce additional end-to-end latency, which demonstrates the effectiveness of the lazy update mechanism in PAT.

### v2

@@ 11. Conclusion
In this work, we present PAT, a prefix-aware attention kernel implementation for LLM decoding. PAT reduces redundant global memory access and improves bandwidth utilization through a pack–forward–merge paradigm. The design incorporates a heuristic pack scheduler, a resource-efficient multi-tile kernel, and forward-stage optimizations that mitigate execution bubbles. We implement PAT as an off-the-shelf plugin to vLLM and demonstrate that it reduces the attention latency by 53.5% on average under synthetic workloads and reduces the TPOT by up to 93.1% under real-world workloads compared with existing works.
12. Acknowledgments
We appreciate the insightful feedback from the anonymous reviewers and our shepherd, Seonjin Na. This work is supported by the National Natural Science Foundation of China under grants No. 62572341 and No. 62202328.
Appendix A Artifact Appendix
A.1. Abstract
We provide the source code of PAT along with scripts to reproduce experimental results presented in the evaluation section. This appendix includes instructions for reproducing the kernel performance evaluation results in Figure 10 and four representative end-to-end performance evaluation results in Figure 11.
To expedite artifact evaluation, a pre-built Docker image is available that contains the fully configured environment, pre-compiled source code, and datasets as specified in §8.2. The experiment require an x86-64 Linux host with at least 64GB RAM, 200GB of free disk space, and an NVIDIA A100 GPU (80GB). For convenience and consistent performance, we recommend using a Google Cloud a2-ultragpu-1g instance with the “Deep Learning VM with CUDA 12.4” system image.
A.2. Artifact check-list (meta-information)
  • Algorithm: Prefix-aware attention kernel implementation.
  • Compilation: Pre-compiled with in a Docker image.
  • Model: Llama3-8B and Qwen3-8B.
  • Run-time environment: Docker, NVIDIA Container Toolkit, and CUDA driver $\geq$ 550.
[…]
  • Metrics: kernel latency, mean TTFT, mean TPOT, P99 TPOT.
  • Output: Figure and console.
  • Experiments: Kernel performance (Figure 10); End-to-end performance (Figure 11).
  • How much disk space required (approximately)?: 100GB.
  • How much time is needed to prepare workflow (approximately)?: 30 minutes.
  • How much time is needed to complete experiments (approximately)?: $\approx$2 hours for kernel performance experiment (Figure 10); $\approx$4 hours for the four selected end-to-end performance experiments (Figure 11).
  • Publicly available?: Yes.
  • Code licenses (if publicly available)?: MIT.
  • Archived (provide DOI)?: 10.5281/zenodo.17943362.
A.3. Description
A.3.1. How to access
A.3.2. Hardware dependencies
Requires an x86-64 Linux host with at least 64GB of RAM, 200GB of free disk space, and an NVIDIA A100 GPU (80GB). For convenience and consistent performance, we recommend using a Google Cloud a2-ultragpu-1g instance with the "Deep Learning VM with CUDA 12.4" system image.
A.3.3. Software dependencies
Docker, NVIDIA Container Toolkit, and NVIDIA driver $\geq$550 are required. Other software has been installed within the provided Docker image.
A.3.4. Data sets
The data sets used in experiments are listed in §8.2 and are pre-downloaded in the Docker image.
A.3.5. Models
The model used in these experiments are listed in §8.2, including Qwen3-8B and Llama-3-8B.
A.4. Installation
  1. (1)
    Clone the GitHub repository.
    \$ git clone
    https://github.com/flashserve/PAT.git
  2. (2)
    Pull the pre-built Docker image.
    \$ # about 50GB, including model weights
    \$ docker pull flashserve/pat:ae
  3. (3)
    Start a Docker container with GPU access and mount the repository.
    \$ docker run -it --gpus all --shm-size=64g \
    -v \${PWD}/PAT:/workspace/PAT \
    -w /workspace/PAT \
    flashserve/pat:ae /bin/bash
A.5. Experiment workflow
  1. (1)
    Run the kernel performance experiments. This experiment takes about 1.5 hours to complete.
    \$ cd /workspace/PAT/benchmark
    \$ bash ./run_kernel_bench.sh
  2. (2)
    Run the end-to-end serving performance experiments. Note that completing all experiments requires over 60 GPU-hours, so we provide two scripts for convenience: (1) run_e2e_bench_part.sh: runs a subset of experiments (QPS=7, all workloads, all baselines) for quick verification; (2) run_e2e_bench_full.sh: runs all experiments in Figure 11.
    \$ cd /workspace/PAT/benchmark
    \$ # Quick verification (4-5 GPU-hours)
    \$ bash ./run_e2e_bench_part.sh
    \$ # Full experiments (over 60 GPU-hours)
    \$ # bash ./run_e2e_bench_full.sh
A.6. Evaluation and expected results
  1. (1)
    Generate plots.
    \$ cd /workspace/PAT/plot
    \$ python eval_kernel_perf.py --log-file \
    ../benchmark/kernel_perf.json
    \$ python eval_e2e_from_jsonl.py --log-file \
    ../benchmark/e2e_perf.jsonl
  2. (2)
    Plots similar to Figure 10 and Figure 11 can be found in the fig directory.